Hi could you help me figure out this word problem?
The electric current I in amperes of a circuit is given by the formula log base 2 of I=-t. Find the current when t is 3s.
Thanks for the help

Answer by stanbon(75887)

The electric current I in amperes of a circuit is given by the formula log base 2 of I=-t. Find the current when t is 3s.
--------------------
-t = log(2)I
-3s = log(2)I
I = 2^(-3s)
OR
I = (1/8)^s
================
Cheers,
Stan H.